1.) Two gases have the same temperature but different pressures. The kinetic molecular theory does not predict that

a. molecules in both gases have the same average kinetic energies.

b. molecules in the low-pressure gas travel farther before they collide with other molecules.

c. both gases have the same densities.

d. all collisions of the molecules are elastic.

2.) On the basis of molecular structure and polarity, which of the following has the greatest solubity?

a. CH4

b. CCl4

c. NH3

d. PH3

3.) Which of the following liquid substances has the weakest intermolecular force?

a. C6H6

b. C2H5OH

c. CH3OH

d. C2H6O2

Please provide a short explanation. Thank you!
Answer #1

1) According to molecular theory of gas the collission between gas molecules are elastic , as the energy conservation is followed. Along with it at constant temperature the kinetic energy of all gas molecules under collision is same . The density of gas molecules are also same as they are in constant motion and always collide with each other. So, option a,c,d are correct . But as we know Mean free path which is the average distance covered between two successive collission is given by (lambda) = 1/√2 π Sigma^2 N* we know N* = PNA/RT so, lambda is inversely proportional to pressure. That means at low pressure molecule covered large distance before collission . 2) Both CH4 and CCl4 have tetrahedral geometry having no dipole moment. So, they behave as non polar molecule. This are mainly soluble in non polar solvent . But NH3 and PH3 having pyramidal structure. As, the P and H have almost same electronegativity so it behaves like nonpolar molecule, so it is insoluble in H20 like other two (CH4 and CCl4) . If we considered NH3 then N -H bond is highly polar and it can carry out H bonds with H20 , so it is polar and highly soluble in H20. 3) In B and C both are alcohol having OH group so it can participate easily in intermolecular H bonding . But for other two that is C6H6 and C2H6O2 this type of H bond formation possibility is not present. In C2H6O2 (di methyl peroxide) dipole dipole intermolecular interaction is present while benzene (C6H6) is non polar which possess weak London dispersion force. So, weakest intermolecular force is operate on C6H6.